# No Thing, Nothing: Highlighting Safety-Critical Classes for Robust LiDAR Semantic Segmentation in Adverse Weather

### v1

@@ 1 Introduction
This difference arises because things classes are generally small and dispersed, leaving them vulnerable to missing points that can distort their structure. For example, person class, belonging to the things category, have dense and clear shapes under clean weather, as shown on the far left of the figure. Consequently, under adverse weather conditions, the person and car points are fewer in number and have obscure shapes. We analyze that weather-induced corruption blurs the boundary between things and stuff, leading to frequent misclassification. Even the state-of-the-art method, SJ+LPD [20], often confused between things and stuff as shown in Fig. 2 (b).
@@ 2.2 LiDAR under Adverse Weather Conditions
Recent task-agnostic approaches [14, 28, 10] employ general machine learning techniques, such as teacher-student frameworks and prototyping, to enhance robustness. Kong et al. [14] introduced random dropout in the teacher-student architecture to build noise-resistant models. PointDR [28] generates a prototype in the clean weather branch and multiplies corrupted branch features by this prototype matrix, utilizing clean data insights. DGUIL [10] addresses point uncertainty under domain shifts by by improving point features across varying distributions in uncertain point clouds.
@@ 3.3 Feature Binding Loss
Building on the analysis in Sec. 3.1, we first introduce a concept of Feature Binding loss (FB). The purpose of FB loss is to “bind” the output features of LSS model, preventing misprediction of things objects as stuff when local corruption severely affects semantic-level features.
To achieve this, we create prototypes for superclasses by grouping visually similar pre-defined classes during LSS training. Tab. S10 summarizes the superclass categories and their corresponding classes. Six superclasses are defined for SemanticKITTI [3], while four superclasses are defined for SemanticPOSS [19]. Following the method in  [12], we associated each class with its respective superclass. Classes with significantly different shapes, such as trunk and vegetation are assigned to separate superclasses. Additionally, classes previously treated as background, like a motorcyclist, were reassigned to the person class.
@@ 4.2 Main Results
SemanticPOSS to SemanticSTF. As shown in Table 3, our method establishes a new state-of-the-art on the SemanticPOSS-to-SemanticSTF benchmark, surpassing the performance of the previous model, SJ+LPD [20], by an impressive +7.9 mIoU. Importantly, in the person category, our method achieves a +15.5 mIoU gain over SJ+LPD. These results underscore the effectiveness of our approach in providing highly accurate segmentation signals for things classes under challenging adverse conditions. The SemanticPOSS dataset contains a higher proportion of things classes at 28.9%, compared to SemanticSTF, and is known for its diversity. Consequently, the model has already learned relatively rich features, resulting in less performance improvement for things classes compared to SemanticKITTI.
@@ 4.3 Ablataion Study
Feature Binding Loss. We conduct an ablation study to evaluate the impact of various prototype strategies of Feature Binding. We compare three approaches: assigning prototypes to each class individually, grouping semantically similar classes into superclasses, and broadly categorizing classes into coarse prototypes as stuff or things. As shown in Table 4, using superclass prototypes improves things classes performance by +3.3 mIoU compared to the class-wise method. This improvement indicates that our FB method effectively bounds things classes, which are prone to be misclassified as stuff, within their respective superclasses. The coarse prototype approach performs well for things classes but reduces stuff class performance due to the use of a single prototype for numerous samples. In contrast, the superclass prototype strategy achieves balanced results across both things and stuff, providing a +3.5 mIoU boost in things classes. This shows that creating prototypes by grouping classes with high visual similarity into superclasses is more effective for feature binding. Although the coarse prototype method enhances things classes, it causes a 1.5 mIoU decrease in stuff and a 0.2 mIoU drop in overall mIoU compared to the superclass prototype approach. This result shows that combining classes with low visual similarity into a single prototype provides suboptimal gains to things classes. Besides, improved prediction of things in driving scenes is beneficial, but poor prediction of stuff can disrupt long-term planning. Thus, the coarse prototype approach cannot be considered to produce ideal results.
@@ 5 Conclusion
  • Sec. S11 analyzes the mIoU drop for the bicyclist class.

### v2

@@ 1 Introduction
This difference arises because things classes are generally small and dispersed, leaving them vulnerable to missing points that can distort their structure. For example, person class, belonging to the things category, have dense and clear shapes under clean weather, as shown on the far left of the Fig. 2 (a) and (b). Consequently, under adverse weather conditions, the person and car points are fewer in number and have obscure shapes. We analyze that weather-induced corruption blurs the boundary between things and stuff, leading to frequent misclassification. Even the state-of-the-art method, SJ+LPD [20], often confused between things and stuff as shown in Fig. 2 (b).
@@ 2.2 LiDAR under Adverse Weather Conditions
Recent task-agnostic approaches [14, 28, 10] employ general machine learning techniques, such as teacher-student frameworks and prototyping, to enhance robustness. Kong et al. [14] introduced random dropout in the teacher-student architecture to build noise-resistant models. PointDR [28] generates a prototype in the clean weather branch and multiplies corrupted branch features by this prototype matrix, utilizing clean data insights. DGUIL [10] addresses point uncertainty under domain shifts by improving point features across varying distributions in uncertain point clouds.
@@ 3.3 Feature Binding Loss
Building on the analysis in Sec. 3.1, we first introduce a concept of Feature Binding loss (FB). The purpose of FB loss is to “bind” the output features of LSS model, preventing misprediction of things objects as stuff when weather corruption severely affects semantic-level features.
To achieve this, we create prototypes for superclasses by grouping visually similar pre-defined classes during LSS training. Tab. 1 summarizes the superclass categories and their corresponding classes. Six superclasses are defined for SemanticKITTI [3], while four superclasses are defined for SemanticPOSS [19]. Following the method in  [12], we associated each class with its respective superclass. Classes with significantly different shapes, such as trunk and vegetation are assigned to separate superclasses. Additionally, classes previously treated as background, like a motorcyclist, were reassigned to the person class.
@@ 4.2 Main Results
SemanticPOSS to SemanticSTF. As shown in Table 3, our method establishes a new state-of-the-art on the SemanticPOSS-to-SemanticSTF benchmark, surpassing the performance of the previous model, SJ+LPD [20], by an impressive +7.9 mIoU. Importantly, in the person category, our method achieves a +15.5 mIoU gain over SJ+LPD. These results underscore the effectiveness of our approach in providing highly accurate segmentation signals for things classes under challenging adverse conditions. The SemanticPOSS dataset contains a higher proportion of things classes at 28.9%, compared to SemanticSTF, and is known for its diversity. Consequently, the model has already learned relatively rich features, resulting in less performance improvement for things classes compared to SemanticKITTI-to-SemanticSTF benchmark.
@@ 4.3 Ablataion Study
Feature Binding Loss. We conduct an ablation study to evaluate the impact of various prototype strategies of Feature Binding. We compare three approaches: assigning prototypes to each class individually, grouping semantically similar classes into superclasses, and broadly categorizing classes into coarse prototypes as stuff or things. As shown in Table 4, using superclass prototypes improves things classes performance by +3.3 mIoU compared to the class-wise method. This improvement indicates that our FB method effectively bounds things classes, which are prone to be misclassified as stuff, within their respective superclasses. The coarse prototype approach performs well for things classes but reduces stuff class performance due to the use of a single prototype for numerous samples. In contrast, the superclass prototype strategy achieves balanced results across both things and stuff, providing a +3.5 mIoU boost in things classes. This shows that creating prototypes by grouping classes with high visual similarity into superclasses is more effective for feature binding. Although the coarse prototype method enhances things classes, it causes a 1.5 mIoU decrease in stuff and a 0.2 mIoU drop in overall mIoU compared to the superclass prototype approach. This result shows that combining classes with low visual similarity into a single prototype provides suboptimal gains. Besides, improved prediction of things in driving scenes is beneficial, but poor prediction of stuff can disrupt long-term planning. Thus, the coarse prototype approach cannot be considered to produce ideal results.
@@ 5 Conclusion
  • Sec. S11 analyzes the mIoU drop for the bicyclist and fence class.
  • Sec. S12. provides discussions about the performance degradation of car class on SemanticPOSS-to-SemanticSTF benchmark.
  • Sec. S13. provides related works on subclass- or prototype- based methods.
  • Sec. S14. provides additional experiments on different weather simulations.
  • Sec. S15. provides additional experiments on various datasets.
  • Sec. S16. provides additional experiments for different superclass criterions.